Project

General

Profile

Actions

2017-09-14-LH2

Phone details

  1. Actions
    • AK/PW: test permissions of required connections HMI to PLC.
      • tbc-CW
    • PW/MT: to test PV20 in sequence.
      • stands until confirmed by MT?
    • AK: to arrange meeting on archiver with JB and implement. JB to confirm
    • AG: Check bonding of lightning conductor - with Richard Hale CW to pursue.
    • SB: create documentation necessary to support shifter training. complete.
  2. Monitoring and archival of data - PF/AK
    • ISIS MCR GUI working.
    • Audio tbc with Ajit - JB
  3. Current technical
    • helium dewar - MT to order for next week.
    • vacuum pumps - move plan. AN.
  4. Documentation
    • Sign-off list. Complete.
    • Modified induction training. - complete.
      DC instructions posted.
    • if H2 detectors alarms go off LH2 expert intervention?
      LH2 group to detail response if hydrogen detectors in hall are triggered.
  5. Timetable
    • Cooldown.
      • Set temperature control at 19K to 20K
      • wait until condenser hits 19K, allow to regulate for a few hours- possibly through Sat/Sat night.
      • Switch to H2 on Sunday morning.
        • babysitters from Sunday evening. 4pm-00, 00-8am
        • Ajit to close expert GUI.
  6. Expert cover
    • Rota to be posted with contact numbers.
  1. Next meeting
  2. AoB

Files:

Updated by Whyte, Colin over 5 years ago · 2 revisions